# Quantum-Classical Hybrid Quantized Neural Network

### v1

@@ Abstract
In this paper, we present a novel Quadratic Binary Optimization (QBO) model for quantized neural network training, enabling the use of arbitrary activation and loss functions through spline interpolation. We introduce Forward Interval Propagation (FIP), a method designed to tackle the challenges of non-linearity and the multi-layer composite structure in neural networks by discretizing activation functions into linear subintervals. This approach preserves the universal approximation properties of neural networks while allowing complex nonlinear functions to be optimized using quantum computers, thus broadening their applicability in artificial intelligence. We provide a theoretical upper bound on the number of ising spins required, which is independent of the training sample size. A significant challenge in solving the associated Quadratic Constrained Binary Optimization (QCBO) model is the presence of numerous constraints. When employing the penalty method to handle these constraints, tuning a large number of penalty coefficients becomes a critical hyperparameter optimization problem, increasing computational complexity and potentially affecting solution quality. To address this, we employ the Quantum Conditional Gradient Descent (QCGD) algorithm, which leverages quantum computing to directly solve the QCBO problem. We prove the convergence of QCGD under a quantum oracle with randomness and bounded variance in objective value, as well as under limited precision constraints in the coefficient matrix. Additionally, we provide an upper bound on the Time-To-Solution for the QCBO solving process. Experimental results using a coherent Ising machine (CIM) demonstrate a $94.95\%$ accuracy on the Fashion MNIST classification task, with comparisons to classical algorithms highlighting superior efficiency. We further validate the robustness of the spline interpolation method and analyze the impact of varying precision levels on QCGD convergence, underscoring the reliability of our framework.
1 Introduction
@@
The rapid advancement of deep learning has transformed a wide array of applications, from natural language processing [1] to computer vision [2], enabling intelligent systems to permeate our daily lives through devices such as smartphones [3] and autonomous vehicles [4]. However, the computational complexity and resource demands of deep neural networks (DNNs), typically trained and executed in high-precision floating-point arithmetic (e.g., 32-bit FP32), pose significant challenges for deployment in resource-constrained environments [5]. Neural network quantization has emerged as a pivotal technique to address these challenges by converting high-precision floating-point weights and activations into lower-precision formats, reducing memory usage, computational demands, and energy consumption with minimal accuracy loss [6]. This technique enables DNNs to operate effectively on resource-limited edge devices such as smartphones [7].
@@
Training quantized neural networks presents unique challenges despite their efficiency in inference. Unlike conventional deep neural networks, which rely on gradient-based optimization in a continuous domain, quantized models require solving a discrete optimization problem. The quantization process restricts weights and activations to a predefined set of discrete values, resulting in a combinatorial search space and landscape [8]. Research on quantized neural network training has primarily focused on two approaches, Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT) [6]. PTQ transforms a pre-trained full-precision model into a low-bit representation by applying techniques such as range calibration or adaptive rounding, exemplified by AdaRound [9]. This method is computationally efficient but tends to suffer from significant accuracy loss under aggressive quantization, particularly at 4-bit precision or lower. In contrast, QAT integrates quantization effects during training, leveraging simulated quantization operations and the Straight-Through Estimator (STE) to improve accuracy in low-precision scenarios [10]. However, despite its advantages, QAT relies on heuristic approximations to circumvent the challenges posed by discrete optimization, limiting its ability to fully exploit the combinatorial structure of the problem.
@@
These limitations highlight the need for a fundamentally different approach to training quantized neural networks—one capable of efficiently navigating the discrete optimization landscape. Quantum computing offers a promising avenue to address this challenge. Delving deeper into this convergence, the role of coherent Ising machines (CIM) [11, 12, 13], becomes a focal point of exploration. Leveraging the principles of quantum coherence and Ising spin dynamics, these systems offer a unique approach to efficiently solving optimization problems central to neural network training, as well as in other fields such as wireless communication [14, 15], molecular docking [16], computer vision [17, 18, 19, 20, 21, 22], etc.
In the realm of neural network training, the choice of activation function plays a crucial role in determining the network’s performance and adaptability to different tasks [23]. While certain problems may benefit from the simplicity and efficiency of ReLU, others may require the nuanced behavior of Sigmoid or Tanh functions. Restricting the network to a specific activation function limits its expressive power and adaptability, potentially leading to suboptimal performance on certain tasks. Therefore, developing training methods that can accommodate any activation function is paramount, as it enables neural networks to flexibly adapt to a wide range of problems and datasets, enhancing their practicality and utility. A pioneering effort in this direction is the Completely Quantum Neural Network (CQNN) framework [24], which embeds an entire neural network—including weights, biases, activation functions, and loss minimization—into a quantum annealer, achieving training in a single annealing step. This approach demonstrates advantages such as consistent convergence to the global minimum and reduced training time, contrasting with the iterative nature of classical gradient descent [24].
@@
However, a critical limitation of CQNN emerges in its handling of activation functions. To conform to the quadratic constraints of the Ising model Hamiltonian, CQNN approximates arbitrary activation functions (e.g., ReLU, sigmoid) with polynomials [24]. This choice raises concerns rooted in approximation theory. As established by Leshno et al. [25], multilayer feedforward networks with non-polynomial activation functions are dense in the space of continuous functions between Euclidean spaces under the compact convergence topology, a property ensuring universal approximation capability. Polynomial activation functions, conversely, limit this expressiveness, as they form a finite-degree polynomial space incapable of approximating arbitrary continuous functions [25, 26]. In addition to the limitations in expressive power, another significant drawback arises when constructing QUBO models is, polynomial activation functions require a large number of ising spins for degree reduction. This issue further exacerbates the challenges of applying CQNN in practical scenarios. Therefore, adapting arbitrary, non-polynomial activation functions to quantum computing frameworks remains an open challenge, requiring innovative encoding strategies or hybrid quantum-classical solutions to preserve both computational efficiency and model versatility.
@@
When addressing constrained quadratic optimization problems, the penalty method is frequently used to incorporate constraints into the objective function [27]. However, this approach has significant limitations. First, quadratic penalty functions introduce pairwise interactions among all variables, transforming sparse problems into fully connected ones. This increased connectivity requires complex mapping techniques, particularly for quantum annealers with limited connectivity [28], substantially raising resource demands on near-term quantum hardware [29]. Second, quadratic penalties generate large energy scales that can exceed hardware dynamic ranges, necessitating Hamiltonian normalization. This normalization reduces the problem’s effective energy resolution, impairing optimization performance. Additionally, the penalty method alters the optimization landscape, requiring careful tuning of penalty coefficients, a critical hyperparameter for effective constraint handling. As the number of constraints grows, the search space for optimal penalty coefficients expands exponentially, posing a significant challenge to achieving robust optimization outcomes.
In this paper, we propose a novel framework for training quantized neural networks using CIM. We begin by formulating the quantization training problem as a QUBO problem, then leverage quantum algorithms to overcome the shortcomings of existing methods of training quantized neural network. Our approach aims to harness the computational power of quantum systems to achieve high-accuracy, low-bit quantized models suitable for resource-constrained edge devices, paving the way for the next generation of efficient deep learning systems. The contributions of this paper are summarized as following.
The contributions of this paper are summarized as following.
  • We introduce the first QUBO model that supports arbitrary forms of activation and loss functions, while preserving the universal approximation properties of neural networks. This is achieved through the use of spline interpolation, which enables the integration of complex nonlinear activation and loss functions with quantum computers, thereby expanding the applicability of quantum computing in neural network training. Additionally, we propose Forward Interval Propagation (FIP), a novel method that discretizes activation functions into linear subintervals to effectively manage the non-linearity and multi-layer composite relationships. Furthermore, based on existing results regarding the expressive power of neural networks with spline-interpolated activation functions, we theoretically establish an upper bound on the number of ising spins required for the QUBO model, independent of the number of training samples.
  • Given the numerous constraints inherent in QCBO models for quantized neural network training, we leverage recent advancements in hybrid algorithms designed to solve constrained optimization problems directly. These hybrid algorithms combine classical and quantum computing capabilities to address the challenges posed by constrained optimization. We provide a rigorous proof of convergence for the hybrid algorithm, taking the randomness and error of the quantum oracle into consideration, demonstrating that the algorithm can achieve error compensation. Additionally, we derive an upper bound on the Time-To-Solution (TTS) for QCBO problems using the TTS of quantum machines for QUBO problems. We also discuss the convergence properties of the algorithm under limited precision conditions on the coefficient matrix. To further enhance efficiency, we propose a lazy implementation strategy to avoid redundant computations.
  • Our experimental evaluation employs practical CIM and further the quantum-classical hybrid method to train quantized neural networks. The results demonstrate superior performance metrics on standard image recognition datasets, validating the effectiveness of our approach. We further investigate the impact of rounding matrix coefficients to varying degrees of precision, and verify the algorithm’s convergence under these conditions. The results show that the algorithm converges reliably across all tested precision levels, with the objective gap and constraint residual gap exhibiting similar trends. Finally, we also validate the robustness of the spline interpolation method.
Figure 1 illustrates the workflow of our proposed quantized neural network training framework, using the hybrid approach and CIM. The process begins with data preparation and progresses through the formulation of the loss landscape and solution space of a hypercube structure. Key steps include constraint linearization, the transformation of the problem into a QCBO model, and the iterative solution process using an Ising machine. At each iteration, variables are updated based on the QUBO matrix, ensuring convergence toward the optimal solution. Our results showcase the enormous potential of quantum computing in real-world applications, laying a solid foundation for the use of quantum computing in quantized neural network training.
Figure 1: Flowchart of the Quantized Neural Network Training Process Using CIM and Hybrid Techniques.
1.1 Related Work
The landscape of quantum machine learning research reveals a diverse range of approaches integrating quantum computing with neural network training. The majority of studies operate within the framework of gate quantum computing, while a limited but notable portion explores the adiabatic quantum computing (AQC) paradigm.
Departing from conventional generative training with Deep Belief Networks (DBN), [30] employs quantum annealing machines for sampling and inference, aiming to alleviate the time-consuming nature of generative training. As a related study of our work, [31] explores Binary Neural Networks (BNN), presenting a quadratic unconstrained binary optimization formulation for the training problem. Moving to [24], it delves into embedding and training general neural networks in a quantum annealer without introducing classical elements. This is achieved through polynomial approximation of activation functions and the reduction of binary higher-order polynomials into quadratic ones. However, a drawback is noted in the need for significant additional ising spins when high precision in fitting is required, as the polynomial degree may be quite high, well beyond the second degree, and reducing the order would demand a substantial increase in additional bits. Additionally, [32] investigates the training of multi-layer feedforward networks on Ising machines, establishing specific QUBO models for particular activation functions such as Sign and ReLU. Lastly, [33] introduces a stochastic optimizer tailored for training binary neural networks using real quantum hardware. The paper demonstrates that projecting weights onto binary variables achieves convergence, establishing an equivalence between binary projection and a specific Quadratic Unconstrained Binary Optimization (QUBO) problem that can be implemented on quantum hardware.
2 Preliminaries and Problem Definition
@@
We conceptualize a quantized neural network (QNN) as an abstract parameterized operator $\mathcal{Q}:\mathbb{R}^{d}\to\mathbb{R}$, transforming an input space $\mathbb{R}^{d}$ into an output space $\mathbb{R}$ through a parameter set $\mathcal{W}$ drawn from a discrete parameter space. This operator emerges from a hierarchical cascade of quantized transformations, reflecting the network’s layered and constrained architecture.
@@
The global operator $\mathcal{Q}(x,\mathcal{W})$ is expressed as a nested sequence of $L$ quantized layers:
@@
where $L$ denotes the depth of the network, and $\mathcal{W}$ encompasses the quantized parameters across all layers. For each layer $l$ ($1\leq i\leq L$), the parameter set $\mathcal{W}_{l}=(\mathbf{W}_{l},\mathbf{b}_{l})$ consists of a quantized weight matrix $\mathbf{W}_{l}$ and a quantized bias vector $\mathbf{b}_{l}$, both confined to a finite set of discrete values defined by a quantization scheme (e.g., integer or fixed-point representations), with dimensional compatibility where the input dimension at layer 0 is $d$ and the output dimension at layer $L$ is 1.
Each layer’s transformation $\mathcal{Q}_{l}$ operates on the quantized output of the previous layer via:
@@
where $\sigma:\mathbb{R}\to\mathbb{R}$ is a non-linear activation function applied elementwise, and the operation $\cdot$ reflects a quantized linear combination.
@@
Here, the sum is taken over all the neurons in the $(l-1)$-th layer.
In this paper, we adopt piecewise linear functions as the activation functions within the neural network, as shown in Figure 2. These functions are particularly advantageous due to their simplicity and computational efficiency, while still maintaining the ability to approximate complex, non-linear relationships, as shown in the following fact.
@@
This fact confirms that quantized piecewise linear neural network, regardless of the specific activation, approximate target functions with efficiency comparable to unquantized one, incurring only a low complexity overhead. Ding et al. [34] further reveal a theoretical optimal number of discrete weight values, typically lies between 1 and 4 bits across a wide range of $d$ and $\epsilon$, indicating that effective quantization of piecewise linear neural network requires only a small number of discrete weights. This aligns with empirical evidences in [5, 36]. Such efficiency supports the use of quantized piecewise linear neural network in quantum computing contexts, where minimal discrete states could leverage quantum parallelism for training.
Figure 2: Feedforward Neural Network With Piecewise Linear Activation Function
While Fact 1 establishes the universal approximation capability of quantized piecewise linear neural networks, it remains an interesting question to quantify the approximation error introduced when replacing a smooth activation function with a piecewise linear one, especially in the presence of spectral normalization, which is often employed to control model complexity and improve generalizability [37, 38]. We present the following theorem to bound the error induced by piecewise linear approximations in a feedforward neural network.
@@
Consider an $L$-layer feedforward neural network $f:\mathbb{R}^{d}\to\mathbb{R}^{m}$, where $W_{l}\in\mathbb{R}^{m_{l}\times m_{l-1}}$ represents the weight matrix in the $l$-th layer, and is spectrally normalized [37, 38]. $\sigma:\mathbb{R}\to\mathbb{R}$ is Lipschitz continuous with constant $L_{\sigma}\leq 1$. Let $\hat{f}$ denote the network with $\sigma$ replaced by a piecewise linear function $\hat{\sigma}$, satisfying:
@@
The proof is provided in Appendix B. This result and analysis provides insight into the trade-offs between model efficiency and approximation accuracy, particularly relevant in the context of quantum enhanced neural network training.
3 Dealing Non-linearity With Spline Fitting
Formally we consider the following problem:
$\min_{\mathbf{x}\in\{0,1\}^{n}}f(h(\mathbf{x})),$ (8)
    $S(h(\mathbf{x}))=\sum_{i=1}^{n}S_{i}(h(\mathbf{x}))\cdot\beta_{i},$ (9)
3.1 Piecewise Constant Segment
$\min$ $\sum_{i=0}^{n}{{\beta_{i}}\cdot S(M_{i})}$ (10)
$\sum_{i=1}^{n}{\beta_{i}}=1,\beta_{i}\in\{0,1\}$ (12)
Dealing with the inequality constraint.
We have now derived a QUBO model capable of approximately solving optimization problems with arbitrary objective functions in (3). However, it is important to note that in representing $\beta_{i}$, we introduced two inequality constraints. Here, we introduce the following two approaches:
  • For $s\in[-\frac{1}{2},\frac{1}{2}]$, adding a penalty term
Theorem 2
Example 1 (Sign function)
$\text{sign}(x)$ ($1$ if $x\geq 0$ else $-1$) can be represented as
where $M=\sup|x|$ and $s\in[-\frac{1}{2},\frac{1}{2}]$ is a slack variable.
4 QCBO Model of Neural Network Training
The training problem is abstracted as an optimization over the discrete parameter space to minimize a generalized loss across all inputs $x$:
@@ 4 QCBO Model of Neural Network Training
where $a^{0}_{x}=x$ serves as the input, $a^{L}_{x}$ is the quantized network output, $y_{x}$ is the target output, and $\mathcal{W}_{\text{quant}}$ denotes the set of all permissible quantized parameter configurations. This formulation abstracts the training of a quantized neural network as a discrete optimization problem over a constrained parameter manifold, capturing the trade-offs between accuracy and computational efficiency inherent in quantization.
Activation functions play a crucial role in neural networks by introducing nonlinearities that enable complex mappings between inputs and outputs. In this section, we leverage spline approximation with piecewise constant functions to approximate activation and loss functions, such as mean squared error (MSE), cross-entropy loss, and Huber loss, which quantify the difference between predicted and actual values in machine learning tasks. To establish the QUBO model, we further propose the use of Forward Interval Propagation.
@@
This is designed to address the challenges of non-linearity in neural networks, particularly the complex multi-layer composite relationships inherent in activation functions. The motivation behind FIP is to simplify these complex interactions by discretizing the activation function into multiple linear subintervals. This allows for efficient computation while preserving the expressive power of neural networks.
In FIP, we divide the entire value range into multiple subintervals. The activation function is defined as a piecewise constant function, where each neuron’s output is determined by the specific subinterval it belongs to. The sum of these outputs, along with the bias term, determines the subinterval for the next layer’s output. This process is repeated during forward propagation, establishing a relationship between the input and output by propagating through the defined subintervals.
@@
Figure 3: Schematic representation of Forward Interval Propagation (FIP) in a neural network, illustrating the discretization of activation functions into linear subintervals. The diagram shows multiple input neurons (left) connected through weights $(w_{0},w_{1},\ldots,w_{m})$ and a bias term ($b$) to a summation node ($\sigma$), followed by an activation function. The green highlighted intervals indicate the specific subintervals that the activation function’s output values belong to, determining the output subinterval for the next layer, effectively capturing multi-layer composite relationships during forward propagation.
We divide the interval $[\underline{M},\overline{M}]$ as $\underline{M}=M_{0}\leq M_{1}\leq M_{2}\leq\ldots\leq M_{n}=\overline{M},$ and the neural network training problem is formulated as the following QCBO problem:
@@
$\sum_{i=1}^{n}{\beta^{(i-1)}_{z^{l}_{x,j}}}M_{i-1}\leq\sum_{k}{w^{l}_{jk}a^{l-1}_{x,k}+b^{l}_{j}}\leq\sum_{i=1}^{n}{\beta^{(i)}_{z^{l}_{x,j}}}M_{i},\forall x,j,l$ (19)

### v2

@@ Abstract
Here in this work, we present a novel Quadratic Binary Optimization (QBO) model for quantized neural network training, enabling the use of arbitrary activation and loss functions through spline interpolation. We introduce Forward Interval Propagation (FIP), a method designed to tackle the challenges of non-linearity and the multi-layer composite structure in neural networks by discretizing activation functions into linear subintervals. This approach preserves the universal approximation properties of neural networks while allowing complex nonlinear functions to be optimized using quantum computers, thus broadening their applicability in artificial intelligence. We provide theoretical upper bounds on the approximation error and the number of Ising spins required, by deriving the sample complexity of the empirical risk minimization (ERM) problem, from an optimization perspective. A significant challenge in solving the associated Quadratic Constrained Binary Optimization (QCBO) model on a large scale is the presence of numerous constraints. When employing the penalty method to handle these constraints, tuning a large number of penalty coefficients becomes a critical hyperparameter optimization problem, increasing computational complexity and potentially affecting solution quality. To address this, we employ the Quantum Conditional Gradient Descent (QCGD) algorithm, which leverages quantum computing to directly solve the QCBO problem. We prove the convergence of QCGD under a quantum oracle with randomness and bounded variance in objective value, as well as under limited precision constraints in the coefficient matrix. Additionally, we provide an upper bound on the Time-To-Solution for the QCBO solving process. Experimental results using a coherent Ising machine (CIM) demonstrate a $94.95\%$ accuracy on the Fashion MNIST classification task, with only 1.1-bit precision. Compared to classical quantization-aware training (QAT), post-training quantization (PTQ) methods and full precision model, our approach yields higher accuracy, faster training, demonstrating its overall efficiency. In addition, evaluations against full-precision models show significant improvements in memory efficiency and inference latency, with minimal accuracy degradation. We further validate the robustness of the spline interpolation method and analyze the impact of varying precision levels on QCGD convergence, underscoring the reliability of our framework.
Introduction
@@
In recent years, the rapid progress of deep learning has catalyzed significant advances in a wide array of domains, from natural language processing [1] to computer vision [2], allowing intelligent systems to permeate our daily lives through devices such as smartphones [3] and autonomous vehicles [4]. However, the computational complexity and resource demands of deep neural networks (DNNs), typically trained and executed in high-precision floating-point arithmetic (e.g., 32-bit FP32), pose significant challenges for deployment in resource-constrained environments [5]. Neural network quantization has emerged as a pivotal technique to address these challenges by converting high-precision floating-point weights and activations into lower-precision formats, reducing memory usage, computational demands, and energy consumption with minimal accuracy loss [6]. This technique enables DNNs to operate effectively on resource-limited edge devices such as smartphones [7].
@@
However, training quantized neural networks presents unique challenges despite their efficiency in inference. Unlike conventional deep neural networks, which rely on gradient-based optimization in a continuous domain, quantized models require solving a discrete optimization problem. The quantization process restricts the weights and activations to a predefined set of discrete values, resulting in a combinatorial search space and landscape [8]. To overcome these issues, studies on quantized neural network training focused primarily on two approaches: post-training quantization (PTQ) and quantization-aware training (QAT) [8]. PTQ transforms a pre-trained full-precision model into a low-bit representation by applying techniques such as range calibration or adaptive rounding, exemplified by AdaRound [9]. This method is computationally efficient but tends to suffer from significant accuracy loss under aggressive quantization, particularly at 4-bit precision or lower. In contrast, QAT integrates quantization effects during training, leveraging simulated quantization operations and the straight-through estimator (STE) to improve accuracy in low-precision scenarios [10]. Despite its advantages, QAT relies on heuristic approximations to circumvent the challenges posed by discrete optimization, limiting its ability to fully exploit the combinatorial structure of the problem.
@@
These limitations highlight the need for a different approach to training quantized neural networks—one capable of efficiently navigating the discrete optimization landscape. The Ising machine [11, 12, 13, 14, 15, 16, 17] offers a promising avenue to address this challenge. By leveraging the energy-minimization properties of Ising models, researchers have applied Ising machines to train neural networks and perform statistical sampling [18]. For example, photonic Ising machines have demonstrated capabilities in low-rank combinatorial optimization and statistical learning tasks [19]; Equilibrium propagation using Ising machines has been utilized as a biologically plausible alternative to backpropagation for training neural networks effectively [20], and sparse Ising machines have been employed to train deep Boltzmann networks [21]. Using the principles of quantum coherence and Ising spin dynamics, these systems offer a unique approach to efficiently solving optimization problems central to neural network training, as well as in other fields such as wireless communication and signal processing [22, 23], molecular docking [24], computer vision [25, 26, 27, 28, 29, 30], etc.
In neural network training, the selection of the activation function critically influences the performance of the model and the adaptability to the task [31]. While the simplicity and computational efficiency of Rectified Linear Units (ReLUs) make them suitable for many applications, certain scenarios demand the distinctive non-linear characteristics of sigmoid or hyperbolic tangent (tanh) functions. Restricting the network to a specific activation function limits its expressive power and adaptability, potentially leading to suboptimal performance on certain tasks. Consequently, the development of training methodologies capable of supporting diverse activation functions becomes essential, as this versatility allows neural networks to effectively address broader problem domains and dataset characteristics, thereby significantly enhancing their practical applicability. A pioneering effort in this domain is the complete quantum neural network (CQNN) framework [32], which implements a comprehensive neural network architecture encompassing weights, biases, activation functions, and loss minimization within a quantum annealing system. This approach achieves the training in a single annealing step, offering notable advantages, including guaranteed convergence to global optima and substantially reduced training durations, in stark contrast to the iterative optimization process characteristic of conventional gradient descent methods.
@@
However, a critical limitation of CQNN emerges in its handling of activation functions. To conform to the quadratic constraints of the Ising model Hamiltonian, CQNN approximates arbitrary activation functions (e.g., ReLU, sigmoid) with polynomials. This choice raises concerns rooted in approximation theory. As established by Leshno et al. [33], multilayer feedforward networks with non-polynomial activation functions are dense in the space of continuous functions between Euclidean spaces under the compact convergence topology, a property ensuring universal approximation capability. Polynomial activation functions, conversely, limit this expressiveness, as they form a finite-degree polynomial space incapable of approximating arbitrary continuous functions [33, 34]. In addition to the limitations in expressive power, another significant drawback arises when constructing QUBO models is, polynomial activation functions require a large number of Ising spins for degree reduction. This issue further exacerbates the challenges of applying CQNN in practical scenarios. Therefore, adapting arbitrary, non-polynomial activation functions to quantum computing frameworks remains an open challenge, requiring innovative encoding strategies or hybrid quantum-classical solutions to preserve both computational efficiency and model versatility.
@@
When addressing constrained quadratic optimization problems, the penalty method is frequently used to incorporate constraints into the objective function [35]. However, this approach has significant limitations. First, quadratic penalty functions introduce pairwise interactions among all variables, transforming sparse problems into fully connected ones. This increased connectivity requires complex mapping techniques, particularly for quantum annealers with limited connectivity [36], substantially raising resource demands on near-term quantum hardware [37]. Second, quadratic penalties generate large energy scales that can exceed hardware dynamic ranges, necessitating Hamiltonian normalization [37]. This normalization reduces the problem’s effective energy resolution, impairing optimization performance. Additionally, the penalty method alters the optimization landscape, requiring careful tuning of penalty coefficients, a critical hyperparameter for effective constraint handling. As the number of constraints grows, the search space for optimal penalty coefficients expands exponentially, posing a significant challenge to achieving robust optimization outcomes.
In this study, we present a framework for training quantized neural networks using CIM. Our approach begins by formulating the quantization training problem as a QUBO problem and then leveraging quantum algorithms to overcome the shortcomings of existing methods of training quantized neural networks. Furthermore, we derive a sample complexity bound from an optimization perspective, effectively constraining the number of spins required in the quadratic binary model. This bound lays a critical foundation for scaling to large-scale scenarios by ensuring computational feasibility and efficiency. Our experimental evaluation highlights the effectiveness of this approach across multiple dimensions. Compared to standard quantization-aware training (QAT) methods, our method achieves higher accuracy with training times reduced by several orders of magnitude. When evaluated against post-training quantization (PTQ) algorithms, it delivers superior accuracy with fewer parameter bits—achieving $94.95\%$ accuracy with only $1.1$-bit precision. Furthermore, compared to full-precision models with continuous weights, our quantized networks significantly reduce memory footprint and inference latency while maintaining competitive accuracy. These results demonstrate the potential of quantum optimization to enable efficient and high-performing deep learning models for edge deployment, and paves the way for the next generation of efficient deep learning systems.
Results
Principle
@@ Results
The approach first conceptualizes a quantized neural network (QNN) as a parameterized operator $\mathcal{Q}:\mathbb{R}^{d}\to\mathbb{R}$, which transforms an input space $\mathbb{R}^{d}$ into the output space $\mathbb{R}$ through a parameter set $\mathcal{W}$ drawn from a discrete parameter space. This operator emerges from a hierarchical cascade of quantized transformations, reflecting the network’s layered and constrained architecture.
@@
The global operator $\mathcal{Q}(x,\mathcal{W})$ could be expressed as a nested sequence of $L$ quantized layers:
@@
where $L$ denotes the depth of the network, and $\mathcal{W}$ encompasses the quantized parameters across all layers. For each layer $l$ ($1\leq i\leq L$), the parameter set $\mathcal{W}_{l}=(\mathbf{W}_{l},\mathbf{b}_{l})$ consists of a quantized weight matrix $\mathbf{W}_{l}$ and a quantized bias vector $\mathbf{b}_{l}$, with dimensional compatibility where the input dimension at layer $1$ is $d$ and the output dimension at layer $L$ is 1.
Each layer’s transformation $\mathcal{Q}_{l}$ operates on the quantized output of the previous layer via the transform as:
@@
where $\sigma:\mathbb{R}\to\mathbb{R}$ denotes the non-linear activation function applied elementwise, and the operation $\cdot$ reflects a quantized linear combination.
@@
Here, the sum of $k$ is taken over all the neurons in the $(l-1)$-th layer.
Figure 1 illustrates the workflow of the quantized neural network training framework, using the hybrid approach and CIM. The process begins with data preparation and progresses through the formulation of the loss landscape and solution space of a hypercube structure. Key steps include the constraint linearization, the transformation of the problem into a QCBO model, and the iterative solution process using an Ising machine. Variables are updated based on the QUBO matrix at each iteration, ensuring convergence toward the optimal solution. Therefore, this work can achieve marked performance in quantum computing in real-world applications, laying a solid foundation for using quantum computing in QNN training.
Figure 1: The flowchart of the quantized neural network training process using CIM and hybrid techniques.
Spline-Induced Neural Network and Optimization Landscape
The piecewise linear functions are adopted as the activation functions in the neural network. These functions are particularly advantageous due to their simplicity and computational efficiency, while still maintaining the ability to approximate complex, non-linear relationships, as shown in the following fact.
@@
This fact confirms that a quantized piecewise linear neural network, regardless of the specific activation, approximates target functions with efficiency comparable to an unquantized one, incurring only a low complexity overhead. Ding et al. [38] further reveal a theoretical optimal number of discrete weight values, typically lies between 1 and 4 bits across a wide range of $d$ and $\epsilon$, indicating that effective quantization of piecewise linear neural network requires only a small number of discrete weights. The efficiency supports the use of quantized piecewise linear neural networks in quantum computing contexts, where minimal discrete states could leverage quantum parallelism for training.
While Fact 1 establishes the universal approximation capability of quantized piecewise linear neural networks, it remains an interesting question to quantify the approximation error introduced when replacing a smooth activation function with a piecewise linear one, especially in the presence of spectral normalization, which is often employed to control model complexity and improve generalizability [40]. In this work, we present the following theorem to provide the error bound induced by piecewise linear approximations in a feedforward neural network.
@@
Consider an $L$-layer feedforward neural network $f:\mathbb{R}^{d}\to\mathbb{R}^{m}$, where $W_{l}\in\mathbb{R}^{m_{l}\times m_{l-1}}$ represents the weight matrix in the $l$-th layer, and is spectrally normalized [40]. $\sigma:\mathbb{R}\to\mathbb{R}$ is Lipschitz continuous with constant $L_{\sigma}\leq 1$. Let $\hat{f}$ denote the network with $\sigma$ replaced by a piecewise linear function $\hat{\sigma}$, satisfying:
@@
The proof of Theorem 1 is provided in the supplementary material. This result and analysis provide insight into the trade-offs between model efficiency and approximation accuracy, particularly relevant in the context of quantum-enhanced neural network training.
The objective of training can be formulated as a discrete optimization problem over the parameter space, aiming to minimize a generalized loss function evaluated across the entire input distribution $x$:
@@
where $a^{0}_{x}=x$ serves as the input, $a^{L}_{x}$ is the quantized network output, $y_{x}$ represents the target output, and $\mathcal{W}_{\text{quant}}$ denotes the set of all permissible quantized parameter configurations. This formulation captures the trade-offs between accuracy and computational efficiency inherent in quantization.
@@
Activation functions play a crucial role in neural networks by introducing nonlinearities that enable complex mappings between inputs and outputs. We leverage spline approximation with piecewise constant functions to approximate activation and loss functions, such as mean squared error (MSE), cross-entropy loss, and Huber loss, which quantify the difference between predicted and actual values in machine learning tasks.
Forward Interval Propagation(FIP) is designed to address the challenges of non-linearity in neural networks, particularly the complex multi-layer composite relationships inherent in activation functions. Also, FIP is to simplify these complex interactions by discretizing the activation function into multiple linear subintervals. This allows for efficient computation while preserving the expressive power of neural networks. As shown in Figure 2, we divide the entire value range into multiple subintervals. The activation function is defined as a piecewise function, where the output of each neuron is determined by the specific subinterval it belongs to. The summation of these outputs, along with the bias term, determines the subinterval for the output of the next layer. This process is repeated during forward propagation, establishing a relationship between the input and output by propagating through the defined subintervals.
@@
Figure 2: (a) Feedforward neural network with piecewise linear activation function. (b) Schematic representation of forward interval propagation (FIP) in a neural network, illustrating the discretization of activation functions into linear subintervals. The diagram shows multiple input neurons (left) connected through weights $(w_{0},w_{1},\ldots,w_{m})$ and a bias term $b$ to a summation node $\Sigma$, followed by an activation function. The green highlighted intervals and values of $\beta_{i}$s indicate the specific subintervals that the activation function’s output values belong to, determining the output subinterval for the next layer, effectively capturing multi-layer composite relationships during forward propagation.
We divide the interval $[\underline{M},\overline{M}]$ as $\underline{M}=M_{0}\leq M_{1}\leq M_{2}\leq\ldots\leq M_{n}=\overline{M},$ And the neural network training problem could be formulated as the following QCBO expression:
@@
$\sum_{i=1}^{n}{\beta^{(i)}_{z^{l}_{x,j}}}M_{i-1}\leq\sum_{k}{w^{l}_{jk}a^{l-1}_{x,k}+b^{l}_{j}}\leq\sum_{i=1}^{n}{\beta^{(i)}_{z^{l}_{x,j}}}M_{i},\forall x,j,l$ (11)